On dissolving 10.8 g glucose in 240 gm of water its boiling point increases by 0.13*C. Calculate the molicular elevation constant of water.Given mol. Mass of glucose=180.

Grade:12

1 Answers

Vikas TU
7 years ago
From delta Tf = i*Kf*m
0.13 = 1*Kf*(10.8/180)/(240/1000) =>  Kf  = 0.13/0.25 => 0.0052
This would be the constant.